OpenStudy (anonymous):

Paige was assigned some math problems for homework. She solved the first 32 of the problems before dinner. She finished the last 20% of assignment after dinner. How many math problems was Paige assigned

OpenStudy (anonymous):

Well, you know that 32 problems are 80 percent of the total, since she completed 20% of the remaining later.

OpenStudy (anonymous):

32 + x = total problems

OpenStudy (anonymous):

x = (.2)*total problems

OpenStudy (anonymous):

So setup your equation like this: (0.2)(32 + x) + 32 = 32 + x that is (20 percent) + (80 percent) = (total)

OpenStudy (anonymous):

and solve for x

OpenStudy (anonymous):

6.4 + 0.2x + 32 = 32 + x 38.4 + 0.2x = 32 + x 6.4 = 0.8x x = 8 and what was x? x was 20% of the problems (because 32 was 80%)

OpenStudy (anonymous):

so then 32 + 8 = 40 problems total

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Thanks to you bro

OpenStudy (anonymous):

where did you get 6.4 from?